OpenStudy (anonymous):

Which of the following equations is the slope-intercept form of 4x - 2y = 8?

OpenStudy (kva1992):

y=mx+b

OpenStudy (anonymous):

slope-intercept formula is y=mx+b subtract x on both sides -2y=-4x+8 now divide both sides by -2 y=4/2-4

OpenStudy (texaschic101):

you forgot the x

OpenStudy (anonymous):

y=4/2x-4